OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = 8/x and g(x)=8/x i don't get what to do when you have 8/8/x ??? some one please help

OpenStudy (anonymous):

\[f \left( g \left( x \right) \right)=\frac{ 8 }{ g \left( x \right) }\] put in g(x) hint: \[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b }\times \frac{ d}{ c }\]

OpenStudy (anonymous):

think of numerator, 8, as \[\frac{ 8 }{ 1 }\]

OpenStudy (anonymous):

ohh right if youre dividing you can multiply by its inverse! thank you

OpenStudy (anonymous):

very good and you're welcome!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

then what do i do with 8/1 x x/8??

OpenStudy (anonymous):

\[\frac{ 8 }{ 1 }\times \frac{ x }{ 8 }=\frac{ 8x }{ 8 }=?\]

OpenStudy (anonymous):

and so, that is the same answer for both and that proves that they are inverses?

OpenStudy (anonymous):

yes. they're inverses, they take you back to where you started. f(x) = g(x) = 8/x, right. so let's say you want to find f(g(3)). that says to plug in the y that you get when you plug in 3 for x in g(x). that y is 8/3 and now you plug that in for x in f(x) and you get 8/(8/3) = (8/1) x (3/8) = 3 that's the original x you started with. it took you back.

OpenStudy (anonymous):

thankiyiou

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

do you know the graph of 1/x? it's symmetric about the line y = x which is why it is its own inverse. Same for y = k/x where k is any non-zero constant